OpenStudy (anonymous):

The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers?

OpenStudy (anonymous):

x + y =110 x - y = 28

OpenStudy (nenadmatematika):

41

hero (hero):

x + y = 110 x - y = 28 2x = 82 x = 41

hero (hero):

This computer is so slow